OpenStudy (anonymous):

The equation below shows the relationship between the temperature in degrees Celsius, C, and degrees Fahrenheit, F: F = 9 over 5C + 32 Which of the following formulas correctly solves for C? C = 5F − 32 over 9 C = 5 F over 9 − 32 C = 9 over 5(F − 32) C = 5 over 9(F − 32)

OpenStudy (anonymous):

F = 9/5 (C + 32) -- multiply both sides by 5/9 5/9 F = C + 32 -- subtract 32 5/9 F - 32 = C -- rearrange C = 5/9 (F - 32)

OpenStudy (anonymous):

Let me explain a little bit... if you multiply both sides by the reciprocal, which is 5/9, then the term on the right side will cancel out 5/9F = 9/5 * 5/9 (C + 32)...which equals 5/9F = (C + 32)....understand ?

OpenStudy (anonymous):

and then you just subtract 32 from both sides to find C